Formatting as Table Styles Office Update To keep up-to-date with security updates, fixes. and improvements, choose Check for Updates. 39 a. day is Rachels 30-birthday Five years ago, Rachel opened a brokerage account when her grand cher gave her S 25,0 O for her 25. birthday Rachel added SZD00。th s account her 26 birthday, $3,000 on her 27h birthday, $4,000 on her 28 birthday, and $5,000 on her 29h birthday. Rachels goal is to have $400,000 in the account by her 40th birthday today, she plans to contribute a fixed amount to the account each year on her birthday. She will make 11 contributions, the first one will occur today, and the final Ⅱ occur on her 40a birthday. Complicating things somewhat s the fct that Rachel plans to withdraw S20 000 from the account on her 3S birthday to finance the down payment home. How large does each of these 1 l contributions have to be for Rachel to reach her goal? Assume the account has eared (nd will continuee eam) an effective return of 12% COVER 10
Answer #1
Interest 12%
Age Year Deposit FV factor Future value
25 0        25,000 5.473566 136,839.14
26 1           2,000 4.887112       9,774.22
27 2           3,000 4.363493     13,090.48
28 3           4,000 3.895976     15,583.90
29 4           5,000 3.47855     17,392.75
30 5 P 3.105848
31 6 P 2.773079
32 7 P 2.475963
33 8 P 2.210681
34 9 P 1.973823
35 10 P-20000 1.762342
36 11 P 1.573519
37 12 P 1.404928
38 13 P 1.2544
39 14 P 1.12
40 15 P 1
192,680.50
Sum of Future value for the contribution made between 25th to 29th 192,680.50
Sum of FV factor for age between 30 to 40 20.6545833
FV of withdrawal of age 35 20000*1.762342
FV of withdrawal of age 35     35,246.84
192680.50-35246.84+ P*20.6546=400000
157433.66+ P*20.6546=400000
P*20.6546=400000-157433.66
P*20.6546= 242566.34
P= 242566.34/20.6546
P=     11,743.94